OpenStudy (anonymous):

factor the trinomial completely 7x^3+14x^2+7x+14

OpenStudy (anonymous):

This is how I do it. First common factor of 7. =7(x^3+2x^2+x+2) =7(x^2+____)(x+______) Because I know that it has to give x^3, so it has to be x^2 * x. Now I just fill in the blanks. How many x^2 do we have? 2. =7(x^2+____)(x+2) How many x's do we have? 1. =7(x^2+1)(x+2) 1*2=2, so that works!